OpenStudy (anonymous):

1. Renting a bicycle at a local Bike shop costs $30 for the first day and $10 for each additional day. Create a chart of values and write an equation to find the total cost, C in terms of the number of days d

OpenStudy (tkhunny):

It needs these features: Price(1 day) = $30 Price(2 days) = $40 = 30 + 10 Price(3 days) = $50 = 30 + 20 = 30 + 2*10 Price(4 days) = $60 = 30 + 30 = 30 + 3*10 Price(5 days) = $70 = 30 + 40 = 30 + 4*10

OpenStudy (anonymous):

No direct answers please, I want liveh elp, not cheatzorz

OpenStudy (tkhunny):

I don't see any direct answers. Let's see your equation.

OpenStudy (anonymous):

Well I got confused pretty much, I need live help, not a "Oh, you need this stuff" then leave, lol

OpenStudy (anonymous):

C = 10d + 30 ______________________________________________________________________ 1| 30 2| 40 3| 50 4| 60 5| 70

OpenStudy (anonymous):

This tkhunny guy isn't very explanatory, so

OpenStudy (anonymous):

@zaibali.qasmi

OpenStudy (anonymous):

R=10d+30

OpenStudy (anonymous):

sorrey.. C=10d+30

OpenStudy (anonymous):

So I did a small chart, I need it checked, and if you can, help me more :D

OpenStudy (anonymous):

Shoot

OpenStudy (anonymous):

Hold on

OpenStudy (anonymous):

C = 10d + 30 ______________________________________________________________________ 1| 30 2| 40 3| 50 4| 60 5| 70

OpenStudy (anonymous):

no.. Best Response Medals 1 C = 10d + 30 ______________________________________________________________________ 1| 30 2| 10*2+30=50 3| 10*3+30=60 4| 70 5| 80 and so..

OpenStudy (anonymous):

Thats confusing :/

OpenStudy (anonymous):

10*2+30=50 is the same as saying 50

OpenStudy (anonymous):

OH WAIT OHHH

OpenStudy (anonymous):

I get it now :)

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

C = 10d + 30 ______________________________________________________________________ 1| 30 2| 10*2+30=50 3| 10*3+30=60 4| 10*4+30=70 5| 10*5+30=80 Right?

OpenStudy (tkhunny):

10d + 30 is incorrect. Try d = 1 10(1) + 30 = 10 + 30 = 40 -- Whoops! Too expensive

OpenStudy (tkhunny):

There are two logical ways to fix it. C = 10(d-1) + 30 C = 10d + 20 Take your pick.

OpenStudy (tkhunny):

What's confusing about "is not correct"? You were doing so well. I want you to show your work. Show some more.

OpenStudy (anonymous):

@SanjanaP

OpenStudy (anonymous):

@tkhunny you realize the problem said 10 for each ADDITIONAL day, not the first day, right? So perhaps the equation does confuse you, but in context, it makes sense.

OpenStudy (tkhunny):

In context, C = 10n + 30 is still incorrect. You mustn't charge the $10 on the first day. Thus, C = 10(n-1) + 30 is correct. Or, after a little algebra, C = 10n + 20.

OpenStudy (anonymous):

@tkhunny @sanjanaP @zaibali.qasmi @DJB8

OpenStudy (tkhunny):

Why are we spamming? This chart is correct. C = 10d + 30 1| 30 2| 40 3| 50 4| 60 5| 70 However, this chart is NOT reproduced by C = 10(n) + 30 Changing the chart is not a good idea. One must change the formula for C.

OpenStudy (anonymous):

So basically you switched the 10d and the 30, and now you're saying you can't change the value of d?

OpenStudy (tkhunny):

No, I'm saying this chart is correct. 1| 30 2| 40 3| 50 4| 60 5| 70 You do not get to change the chart if you write the wrong function. That correct chart is NOT reproduced by C = 10d + 30 or by C = 30 + 10d You must find a function that reproduces the correct chart. You must not alter the correct chart to match an incorrect equation.

OpenStudy (anonymous):

Welp

OpenStudy (tkhunny):

Plenty of stuff on there. You just have to think about it until it soaks in. You'll get it.

OpenStudy (anonymous):

Pfft, I'm slower than a turtle

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@Lyssa123

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

So I need help :/

OpenStudy (anonymous):

hi an ok wat is it -_-

OpenStudy (anonymous):

Is this correct? C = 10d + 30 1| 30 2| 40 3| 50 4| 60 5| 70 When C = total cost, d = the number of days, 30 is the amount to rent a bike for the 1st day and the 10 is the amount to pay for each additional day

OpenStudy (anonymous):

Idk ... this is confusing to me